# Towards Aligning Language Models with Textual Feedback

### v1

@@ Abstract
We present ALT (ALignment with Textual feedback), an approach that aligns language models with user preferences expressed in text. We argue that text offers greater expressiveness, enabling users to provide richer feedback than simple comparative preferences and this richer feedback can lead to more efficient and effective alignment. ALT aligns the model by conditioning its generation on the textual feedback. Our method relies solely on language modeling techniques and requires minimal hyper-parameter tuning, though it still presents the main benefits of RL-based alignment algorithms and can effectively learn from textual feedback. We explore the efficacy and efficiency of textual feedback across different tasks such as toxicity reduction, summarization, and dialog response generation. We find that ALT outperforms PPO for the task of toxicity reduction while being able to match its performance on summarization with only 20% of the samples. We also explore how ALT can be used with feedback provided by an existing LLM where we explore an LLM providing constrained and unconstrained textual feedback. Equal Contribution We also outline future directions to align models with natural language feedback.  [^1] https://github.com/sauc-abadal/ALT
@@
Saüc Abadal Lloret∗1   Shehzaad Dhuliawala∗1   Keerthiram Murugesan2   Mrinmaya Sachan1 1Department of Computer Science, ETH Zürich    2IBM Research {sdhuliawala,sabadal,msachan}@inf.ethz.ch
@@ 1 Introduction
To ensure language models are effective in real-world scenarios, their behavior must be aligned with the specific goals of the applications. Techniques for alignment often involve training a reward model over preference data and using a Reinforcement Learning (RL) solution to steer the model toward preferred responses Ouyang et al. (2022); Snell et al. (2022). A common argument for using RL techniques is that, unlike supervised fine-tuning which trains the model to predict a single good answer, an RL technique allows the model to get both positive and negative reward signals for its predictions (Goldberg, 2023). Reinforcement learning methods, while powerful, often face significant hurdles that hamper their public adoption, i.e., requiring vast amounts of training data Yarats et al. (2021). Vamplew et al. (2022) argue that the scalar reward often provides a very sparse informative signal for the model. color=yellow,size=,fancyline,caption=,]Mrinmaya: What is the main point we are making? RL is suboptimal and inefficient? Or that the existing approach of collecting pairwise preference data and doing RLHF on is expensive and not a natural way for collecting and working with human feedback? As much as I understand, both are linked. However, it seems that textual feedback is our main goal which requires departing from the standard RLHF pipeline and adopting a reward conditional generation approach) and we need to make a clear case for both but in this order? Let me know if you disagree. SD: added more context
This work proposes a text-based feedback mechanism for aligning language models. We propose that providing models with textual feedback, rather than numerical scores, can offer a more nuanced and informative learning signal for understanding human preferences. This textual feedback can improve the process of aligning AI systems. In ALT, we depart from traditional RL approaches such as PPO and focus on reward-conditioned RL. color=yellow,size=,fancyline,caption=,]Mrinmaya: Need a paragraph here summarizing ALT and an argument how it departs from the traditional RLHF approach; why the QUARK style approach allows for efficient modeling of text feedback, and then the experiments after that in a different paragraph. SD: I don’t think Quark is better than PPO and we also find the same… I think Quark + the Text feedback gets it close to PPO. Agreed: sorry, I meant reward conditional RL. Reward-conditioned RL Chen et al. (2021) is an approach that allows the policy to be trained using a supervised learning loss similar to sequence modeling. More recently, reward-conditioned RL has been adapted to the task of alignment (Lu et al., 2022) where they condition generations using reward quantiles as feedback and (Dong et al., 2023b) where the generations are conditioned on numerical reward feedback. Building upon this, our work introduces ALT, which leverages the richness of the signal provided by textual feedback to improve model performance.
@@ 2 ALT: ALignment with Textual feedback
color=yellow,size=,fancyline,caption=,]Mrinmaya: Before jumping into describing your approach, you should discuss and argue why existing RLHF approaches may not be the right thing to do if you want to incorporate textual feedback. Then, you can describe ALT. SD: With traditional RL you are maximizing the expected reward in some way, so there’s no way to incorporate text.
color=yellow,size=,fancyline,caption=,]Mrinmaya: We also need a more higher level description of why ALT works like this? ALT adapts the reward-conditioned RL framework by training the model to be conditioned on textual feedback. ALT (Figure 1) consists of two distinct phases: data collection and training. In the data collection phase, we sample generations from the model and assign language feedback to these generations. In the training phase, the model is trained to map its generations to the assigned feedback. These steps are repeated iteratively as the model is trained to generate conditioned on feedback.
@@ 2.1 Data Collection: Sampling + Feedback
The Sampling in the initial step, uses a supervised fine-tuned model that has not been trained to generate conditioned on feedback. Hence, only for the initial step, we generate conditioned on the input.
@@
After the model has been trained to generate responses conditioned on the feedback, we can now align the model by instructing its generation using text. To generate responses that align the model to certain feedback, exemplar feedback $\hat{f}$ is now used to condition the generation $y_{i}\sim p_{\theta_{k}}(x_{i},\hat{f})$. For example, if the LLM’s generations want to be aligned to be less toxic, feedback Nontoxic can be prepended to the prompt. The intuition behind using exemplar feedback is that we are querying the LLM for its understanding of the feedback we want to steer it toward so we can iteratively refine this representation.
@@ 2.2 Training
$\mathcal{L}_{ref}$ $=\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\sum_{t=1}^{T}\text{KL}\Big{(}p_{0}(y_{i_{t}}|y_{i_{<t}},x_{i})\;||$
$\qquad\qquad p_{\theta}(y_{i_{t}}|y_{i_{<t}},f_{i},x_{i})\Big{)}$
Furthermore, an entropy regularization term to encourage exploration by avoiding too peaky probability distributions can also be added.
$\mathcal{L}_{H}=-\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\text{H}(p_{\theta})$
Finally, the final loss can be written as:
$\mathcal{L}=\mathcal{L}_{LL}+\beta\mathcal{L}_{ref}+\alpha\mathcal{L}_{H}$ (1)
@@
4:Dataset $X$,
5:Datapool $\mathcal{D}$
6:$\mathcal{D}\leftarrow\phi$
@@
21:      $\theta_{k}\leftarrow\operatorname*{arg\,min}_{\theta}\mathcal{L}$
@@ 2.3 Feedback Provider
A pre-existing reward model can be used as a feedback provider by converting its numerical rewards into categorized feedback relevant to the specific task. For instance, in the context of toxicity mitigation, the reward range can be divided into five quantiles, each corresponding to a distinct toxicity level: ‘‘very toxic’’, ‘‘toxic, medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. Our experiments demonstrate that even rule-based textual feedback can enhance model performance compared to using raw scalar reward values or ad-hoc extra embeddings added for each reward quantile. We refer to this variant as $\textsc{ALT}_{\textsc{RM}}$.
LM-based Categorical Feedback
In this approach, we use an existing LLM to generate one of the preset feedbacks. We design a tailored prompt that using in-context learning allows the LM to provide specific textual feedback. Prior work has shown that existing LLMs can be a suitable replacement for custom reward models Lee et al. (2023). We use GPT-3.5-Turbo as a feedback provider. Further details regarding the prompts used for each task can be found in Appendix C. We refer to this variant as $\textsc{ALT}_{\textsc{LMC}}$.
@@ 3.1 Toxicity Reduction
As LLMs are trained on large amounts of text from the internet they can be prone to generating toxic content Gehman et al. (2020). In user-facing settings, we want to steer models away from producing toxic and potentially harmful content. We first investigate the effectiveness of our approach using the RealToxicityPrompts dataset Gehman et al. (2020). Here the LLM is prompted with a seemingly harmless piece of text and is judged on the toxicity of its generation. To assess toxicity, we are focusing on a single generation aspect (the text’s toxicity level), so to get textual feedback we quantize the reward model’s feedback into various preset texts.
3.2 Summarization
We next evaluate if ALT can better help align the LLM to user preferences. To verify this, we experiment using the Reddit TL;DR dataset Völske et al. (2017). The TL;DR dataset has become a common benchmark for measuring alignment. The prompts consist of Reddit forum posts and the task is to generate a summary of the main points of the post while fulfilling different facets that humans care about, such as coherence, accuracy, coverage, or conciseness. Similar to the toxicity task, we leverage an existing reward model that accounts for the different facets humans value on summaries to predict a single scalar reward and quantize the reward model’s feedback into preset texts indicating increasing degrees of alignment fulfillment.
3.3 Dialog Response Generation
For this task, we experiment with the Anthropic HH dataset Bai et al. (2022). The task involves training a model to generate helpful responses to user queries. The model has to learn to balance being helpful without producing content that can cause harm. In this setting, we skip the ad-hoc mapping from quantized numerical scores provided by a reward model to textual feedback and explore using an LLM as the feedback provider by directly providing the textual feedback indicating varying degrees of helpfulness and harmlessness.
3.4 Experimental Details
3.4.1 Toxicity Reduction
We follow the same experimental setup as in Liu et al. (2021); Lu et al. (2022), and consider reducing toxicity from GPT2-large on the RealToxicityPrompts benchmark. In addition, we also conduct an out-of-domain evaluation with the WritingPrompts [^2] https://huggingface.co/datasets/euclaise/writingprompts/viewer/default/test test dataset Fan et al. (2018). As a reward function and a proxy for measuring the toxicity of the LLM generations, we use the Perspective API. [^3] The Perspective API is a service developed by Google that is dynamic and evolves over time. Queries were made from Sep 2023 to Nov 2023. We use $K=5$ quantiles, obtained by sorting the samples in the data pool from lowest toxicity (highest reward) to highest toxicity (lowest reward), and map them to language feedback indicating increasing degrees of toxicity: ‘‘Lowest Toxicity’’, ‘‘Low-Moderate Toxicity’’, ‘‘Moderate Toxicity’’, ‘‘High-Moderate Toxicity’’, and ‘‘Maximum Toxicity’’. We report training details and hyper-parameters in A.1.
@@ 3.4.1 Toxicity Reduction
3.4.2 Summarization
@@
The Reward Model [^5] https://huggingface.co/CarperAI/openai_summarize_tldr_rm_checkpoint used for Quark and $\textsc{ALT}_{\textsc{RM}}$ is a GPT-J model trained on top of the SFT on the TRLX framework using the human preference dataset gathered by Stiennon et al. (2020) for RLHF. We also observed that the $\beta$ on the KL penalty term had little effect on training so we dropped the term in both Quark and $\textsc{ALT}_{\textsc{RM}}$ experiments. We note that we are still able to obtain a lower perplexity than PPO. We report the training details and hyper-parameters in A.2.
@@
3.4.3 Dialog Response Generation
@@ 4 Results
For the task of toxicity reduction (Table 1), we find that $\textsc{ALT}_{\textsc{RM}}$ can reduce the toxicity of the model’s generations more effectively than Quark on both in-domain ($0.148\rightarrow 0.082$), and out-of-domain ($0.193\rightarrow 0.113$) indicating that merely switching out a quantiled reward feedback with textual feedback can result in more effective alignment. We note that ALT outperforms PPO at aligning for lower toxicity while maintaining a lower perplexity ($14.27\rightarrow 12.31$). We provide qualitative examples in B.1.
@@
Figure 3: Training curves for $\textsc{ALT}_{\textsc{LMC}}$ on HH. Evaluation on a held-out validation set. The percentage of Harmless and very helpful generations increases while the percentage of Harmful generations decreases.
@@
Compared to SteerLM, we find that $\textsc{ALT}_{\textsc{LM}}$ is better off at learning to discriminate between ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’, as it almost monotonically increases the % of the former and decreases the % of the latter. We observe that SteerLM fails at improving both the fraction of generations being ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’ compared to the SFT model, and that it becomes less harmful by becoming less helpful, e.g., denying to answer a question or providing an unrelated answer. Regarding ‘‘Harmless and not helpful’’ and ‘‘Harmful’’, both methods follow similar trends but $\textsc{ALT}_{\textsc{LMC}}$ achieves slightly better metrics. We include other training curves showing the generations’ length, percentage of truncated generations, and perplexity in A.3, and a test-time steerability assessment in Appendix E. We provide qualitative results in B.3.
@@
We used a similar set-up as for the TL;DR experiment but we replace the reward model with GPT-3.5-Turbo and prompt it to provide feedback on the summary (Appendix C). We noticed that the summaries produced changed significantly from the SFT model. However, we observed no improvement in the reward model score with the LLM feedback. After 6 iterations we found that the resulting model $\textsc{ALT}_{\textsc{LMU}}$ is comparable to the SFT model and worse than PPO in terms of win rates (Table 2). We hypothesize that the unreliability of GPT-3.5-Turbo in providing unconstrained feedback contributes to this issue. We observed instances where the model gave contradictory feedback on the same summary, potentially hindering a reliable signal for model learning. However, GPT-3.5-Turbo demonstrated more consistency when presented with a constrained set of feedback options, evidenced by $\textsc{ALT}_{\textsc{LMU}}$’s improved performance (Figure 3). We show some examples in B.2.

### v2

@@ Abstract
We present ALT (ALignment with Textual feedback), an approach that aligns language models with user preferences expressed in text. We argue that text offers greater expressiveness, enabling users to provide richer feedback than simple comparative preferences, leading to more efficient and effective alignment. ALT aligns the model by conditioning its generations on the textual feedback. Our method relies solely on language modeling techniques and requires minimal hyper-parameter tuning while retaining the main benefits of RL-based alignment algorithms. We demonstrate the efficacy and efficiency of textual feedback across different tasks, including toxicity reduction, summarization, and dialogue response generation. Notably, ALT outperforms PPO in toxicity reduction and matches its performance on summarization with only 20% of the samples. Equal Contribution We also explore using ALT with feedback from an existing LLM, examining constrained and unconstrained feedback. Additionally, we outline future directions to align models with natural language feedback.  [^1] https://github.com/sauc-abadal/ALT
@@
Saüc Abadal Lloret∗1   Shehzaad Dhuliawala∗1 Keerthiram Murugesan2Mrinmaya Sachan1 1Department of Computer Science, ETH Zürich    2IBM Research {sdhuliawala,sabadal,msachan}@inf.ethz.ch keerthiram.murugesan@ibm.com
@@ 1 Introduction
To ensure language models are effective in real-world scenarios, their behavior must be aligned with the specific goals of the applications. Techniques for alignment often involve training a reward model over preference data and using a Reinforcement Learning (RL) solution to steer the model toward preferred responses Ouyang et al. (2022); Snell et al. (2022). A common argument for using RL approaches is that, unlike supervised fine-tuning which trains the model to predict a single good answer, an RL technique allows the model to get both positive and negative reward signals for its predictions (Goldberg, 2023). Reinforcement learning methods, while powerful, often face significant hurdles that hamper their public adoption, i.e., requiring vast amounts of training data Yarats et al. (2021). Vamplew et al. (2022) argue that the scalar reward often provides a very sparse informative signal for the model.
This work proposes a text-based feedback mechanism for aligning language models. We posit that providing models with textual feedback, rather than numerical scores, can offer a more nuanced and informative learning signal for understanding human preferences. This textual feedback can improve the process of aligning AI systems. In ALT, we depart from traditional RL approaches such as PPO and focus on reward-conditioned RL. Reward-conditioned RL Chen et al. (2021) is an approach that allows the policy to be trained using a supervised learning loss similar to sequence modeling. More recently, reward-conditioned RL has been adapted to the task of alignment in (Lu et al., 2022), where generations are conditioned using reward quantiles as feedback, and in (Dong et al., 2023b), where they are conditioned on numerical reward feedback. Building upon this, our work introduces ALT, which leverages the richness of the signal provided by textual feedback to improve model performance.
@@ 2 ALT: ALignment with Textual feedback
ALT adapts the Decision Transformer (Chen et al., 2021) by training the model to be conditioned on textual feedback thus simplifying the RL setup to conditional supervised-finetuning. Prior works have made use of this framework for alignment, by using reward quantiles (Lu et al., 2022), numerical scores (Dong et al., 2023b), or contrastive feedback (Liu et al., 2023). ALT differs from these approaches by making use of textual feedback.
ALT (Figure 1) consists of two distinct phases: data collection and training. In the data collection phase, we sample generations from the model and assign language feedback to these generations. In the training phase, the model is trained to map its generations to the assigned feedback. These steps are repeated iteratively as the model is trained to generate conditioned on feedback.
@@ 2.1 Data Collection: Sampling + Feedback
In the sampling step, we sample generations from the model conditioned on the feedback. In the first iteration of ALT, we begin with a supervised fine-tuned (SFT) model that has not yet been trained to generate conditioned on feedback and simply sample generations from the SFT model conditioned on the input.
@@
After the model has been trained to generate responses conditioned on the feedback, we can now align the model by instructing its generation using text. To generate responses that align the model to certain feedback, exemplar feedback $\hat{f}$ is now used to condition the generation $y_{i}\sim p_{\theta_{k}}(x_{i},\hat{f})$. For example, if we want the LLM’s generations to be aligned to be less toxic, feedback Nontoxic can be prepended to the prompt. The key intuition behind conditioning on exemplar feedback is to query the LLM for its understanding of the feedback we want to steer it toward so we can iteratively refine this representation.
@@ 2.2 Training
$\mathcal{L}_{ref}=$
$\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\text{KL}\Big{(}p_{0}(y_{i}|x_{i})\;||\;p_{\theta}(y_{i}|x_{i},f_{i})\Big{)}$
We then add an entropy regularization term to encourage exploration.
$\mathcal{L}_{H}=-\mathbb{E}_{(x_{i},y_{i},f_{i})\sim\mathcal{D}}\ \text{H}\Big{(}p_{\theta}(y_{i}|x_{i},f_{i})\Big{)}$
The final loss can be written as:
$\mathcal{L}_{\theta}=\mathcal{L}_{NLL}+\beta\mathcal{L}_{ref}+\alpha\mathcal{L}_{H}$ (1)
@@
4:Dataset $X$ with $Q$ datapoints,
5:Number of iterations $N$
6:Datapool $\mathcal{D}\leftarrow\phi$
@@
21:      $\theta_{k}\leftarrow\operatorname*{arg\,min}_{\theta}\mathcal{L}_{\theta}$
@@ 2.3 Feedback Provider
A pre-existing reward model can be used as a feedback provider by converting its scalar rewards into categorized feedback relevant to the specific task. For instance, in the context of toxicity mitigation, the reward range can be divided into five quantiles, each corresponding to a distinct toxicity level: ‘‘very toxic’’, ‘‘toxic’’, ‘‘medium toxic’’, ‘‘slightly toxic’’, ‘‘nontoxic’’. Our experiments demonstrate that even rule-based textual feedback can enhance model performance compared to using raw scalar reward values or ad-hoc extra embeddings added for each reward quantile. We refer to this variant as $\textsc{ALT}_{\textsc{RM}}$.
LLM-based Categorical Feedback
In this approach, we use an existing LLM to generate one of the preset feedbacks. We design a tailored prompt that using in-context learning allows the LLM to provide specific textual feedback. Prior work has shown that existing LLMs can be a suitable replacement for custom reward models Lee et al. (2023). We use GPT-3.5-Turbo as a feedback provider. Further details regarding the prompts used for each task can be found in Appendix C. We refer to this variant as $\textsc{ALT}_{\textsc{LMC}}$.
@@ 3.1 Toxicity Reduction
As LLMs are trained on large amounts of text from the internet they can be prone to generating toxic content Gehman et al. (2020). In user-facing settings, we want to steer models away from producing toxic and potentially harmful content. Here the LLM is prompted with a seemingly harmless piece of text and is judged on the toxicity of its generation. To assess toxicity, we are focusing on a single generation aspect (the text’s toxicity level), so to get textual feedback we quantize the reward model’s feedback into various preset texts. We experiment using the RealToxicityPrompts benchmark, consisting of 100k prompts created to elicit toxic generations. Our splits amount to 85k, 5k, and 10k for the train, validation and test sets respectively. In addition, we also conduct an out-of-domain (OOD) evaluation with the 15k prompts of the WritingPrompts [^2] https://huggingface.co/datasets/euclaise/writingprompts/viewer/default/test test dataset Fan et al. (2018).
Experimental details.
We follow the same experimental setup as in Liu et al. (2021); Lu et al. (2022), and consider reducing toxicity from GPT2-large. As a reward function and a proxy for measuring the toxicity of the LLM generations, we use the Perspective API. [^3] The Perspective API is a service developed by Google that is dynamic and evolves over time. Queries were made from Sep 2023 to Nov 2023. We use $K=5$ quantiles, obtained by sorting the samples in the data pool from lowest toxicity (highest reward) to highest toxicity (lowest reward), and map them to language feedback indicating increasing degrees of toxicity: ‘‘Lowest Toxicity’’, ‘‘Low-Moderate Toxicity’’, ‘‘Moderate Toxicity’’, ‘‘High-Moderate Toxicity’’, and ‘‘Maximum Toxicity’’. We report training details and hyper-parameters in A.1.
@@
3.2 Summarization
We next evaluate if ALT can better help align the LLM to user preferences. To verify this, we experiment using the Reddit TL;DR dataset Völske et al. (2017). The TL;DR dataset has become a common benchmark for measuring alignment. We used the dataset version hosted in HuggingFace [^4] https://huggingface.co/datasets/CarperAI/openai_summarize_tldr, with splits amounting to 117k, 6.5k, and 6.5k for train, validation, and test. The prompts consist of Reddit forum posts and the task is to generate a summary of the main points of the post while fulfilling different facets that humans care about, such as coherence, accuracy, coverage, or conciseness. Once again, we use an existing reward model that accounts for the different facets humans value on summaries to predict a single scalar reward and quantize the reward model’s feedback into preset texts indicating increasing degrees of alignment fulfillment.
Experimental details.
@@ 3.2 Summarization
The Reward Model [^6] https://huggingface.co/CarperAI/openai_summarize_tldr_rm_checkpoint used for Quark and $\textsc{ALT}_{\textsc{RM}}$ is a GPT-J model trained on top of the SFT on the TRLX framework using the human preference dataset gathered by Stiennon et al. (2020) for RLHF. We observed that the $\beta$ on the KL penalty term had little effect on training so we dropped the term in both Quark and $\textsc{ALT}_{\textsc{RM}}$ experiments. We note that we are still able to obtain a lower perplexity than PPO. We report the training details and hyper-parameters in A.2.
@@
3.3 Dialog Response Generation
For this task, we experiment with the Anthropic HH dataset Bai et al. (2022). The task involves training a model to generate helpful responses to user queries. The model has to learn to balance being helpful without producing content that can cause harm. In this setting, we skip the ad-hoc mapping from quantized numerical scores provided by a reward model to textual feedback, and explore using an LLM as the feedback provider by directly providing the textual feedback indicating varying degrees of helpfulness and harmlessness. We employ the dataset hosted in HuggingFace [^8] https://huggingface.co/datasets/Anthropic/hh-rlhf, comprising 161k and 8.5k examples for the train and test set. We only focus on single-turn responses and for evaluation we select 1000 unique prompts from the test set. We remove duplicate examples from both splits.
Experimental details.
@@ 4 Results
For the task of toxicity reduction (Table 1), we find that $\textsc{ALT}_{\textsc{RM}}$ can reduce the toxicity of the model’s generations more effectively than Quark on both in-domain ($0.148\rightarrow 0.082$), and out-of-domain ($0.193\rightarrow 0.113$), indicating that merely switching out a quantiled reward feedback with textual feedback can result in more effective alignment. We note that ALT outperforms PPO at aligning for lower toxicity while maintaining a lower perplexity ($14.27\rightarrow 12.31$). We provide qualitative examples in B.1.
@@
Figure 3: Training curves for $\textsc{ALT}_{\textsc{LMC}}$ on HH. The percentage of Harmless and very helpful generations increases while the percentage of Harmful generations decreases. Each iteration corresponds to  2k training samples.
@@
Compared to SteerLM, we find that $\textsc{ALT}_{\textsc{LM}}$ is better off at learning to discriminate between ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’, as it almost monotonically increases the % of the former and decreases the % of the latter. We observe that SteerLM fails at improving both the fraction of generations being ‘‘Harmless and very helpful’’ and ‘‘Harmless and helpful’’ compared to the SFT model, and that it becomes less harmful by becoming less helpful, e.g., denying to answer a question or providing an unrelated answer. Regarding ‘‘Harmless and not helpful’’ and ‘‘Harmful’’, both methods follow similar trends but $\textsc{ALT}_{\textsc{LMC}}$ achieves better performance. We include other training curves showing the generations’ length, percentage of truncated generations, and perplexity in A.3, and also perform preliminary tests on $\textsc{ALT}_{\textsc{LMC}}$’s ability to steer the LM in Appendix E. We provide qualitative results in B.3.
@@
We used a similar set-up as for the TL;DR experiment but we replace the reward model with GPT-3.5-Turbo and prompt it to provide feedback on the summary (Appendix C). We noticed that the summaries produced changed significantly from the SFT model. However, we observed no improvement in the reward model score with the LLM feedback. After 6 iterations we found that the resulting model $\textsc{ALT}_{\textsc{LMU}}$ is comparable to the SFT model and worse than PPO in terms of win rates (Table 2). We hypothesize that the unreliability of GPT-3.5-Turbo in providing unconstrained feedback contributes to this issue. We observed instances where the model gave contradictory feedback on the same summary, potentially hindering a reliable signal for model learning. However, GPT-3.5-Turbo demonstrated more consistency when presented with a constrained set of feedback options, evidenced by $\textsc{ALT}_{\textsc{LMC}}$’s improved performance (Figure 3). We show some examples in B.2.